OpenStudy (lena772):

The functions f(x) and g(x) are described below. f(x) = 32x + 8 g(x) = 32x - 9 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (the_fizicx99):

How would you solve this

OpenStudy (lena772):

Idk :(

OpenStudy (skullpatrol):

Do you know what the "slope-intercept" form of a line means?

OpenStudy (skullpatrol):

$$\Huge y=mx+b$$

OpenStudy (lena772):

Yes

OpenStudy (skullpatrol):

The number "b" is the point on the graph that the line intercepts the y-axis.

OpenStudy (lena772):

m=slope

OpenStudy (lena772):

yes ik

OpenStudy (skullpatrol):

$$\Huge y=mx+b$$ $$\Huge f(x) = 32x + 8$$ $$\Huge g(x) = 32x - 9$$

OpenStudy (skullpatrol):

how far has b shifted?

OpenStudy (dan815):

just think about what the constant on the end is doing to every x value you put in, how is changing your y values for every x value

OpenStudy (dan815):

how is it* changing you y values

OpenStudy (dan815):

a graph is just a plot of many points over some "domain"(over some x values in your case)

OpenStudy (dan815):

so think back to tables, when you input x=1,x=2,x=3.. what is happening to your Y

OpenStudy (dan815):

when that -9 is there as opposed to that +8

OpenStudy (skullpatrol):